What is the length of the median from point A in a triangle with the the vertices A(5, 7) * B * (7, - 5) and C(- 5, - 1) 


1
Expert's answer
2021-05-19T08:05:44-0400

A (5, 7)

B (7, - 5)

C (- 5, - 1) 


Let the mid - point of BC be P(x, y)

By mid point formula we have


P(x, y) = "\dfrac{x1 + x2}{2}" , "\dfrac{y1 + y2}{2}"


where x1 , y1 are the coordinates of B and x2, y2 are the coordinates of C .

P(x, y) = (1, -3)



Now the length the median from a can be found by calculating the distance between A and P


AP = "\sqrt{(1-5)^2 + (-3 - 7)^2}"


AP = "\sqrt{(-4)^2 + (-10)^2}"


AP = "\sqrt{116}"

AP = 10.7 units.